# LAG: LLM agents for Leaderboard Auto Generation on Demanding

### v1

@@ Appendix C Cost Analysis
We calculate the average number of input & output tokens required to generate a 20-entry leaderboard, along with the cost analysis using different LLMs, as shown in Table 6. The computational cost of all models remains within 14\$, indicating that LAG is also economically efficient. Overall, the LAG framework consumes more input tokens, while the output tokens represent only a small proportion. OpenAI prices output tokens significantly higher than input tokens, often reaching 4-5 times the cost of input tokens. However, considering the disparity in token numbers, the overall cost remains acceptable.

### v2

@@ Appendix G Cost Analysis
We calculate the average number of input & output tokens required to generate a 20-entry leaderboard, along with the cost analysis using different LLMs, as shown in Table 11. The computational cost of all models remains within 14\$, indicating that League is also economically efficient. Overall, the League framework consumes more input tokens, while the output tokens represent only a small proportion. OpenAI prices output tokens significantly higher than input tokens, often reaching 4-5 times the cost of input tokens. However, considering the disparity in token numbers, the overall cost remains acceptable.
Table 11: Cost of API. The open-source Qwen2.5-7/14B model is deployed on a local server comprising four A800 GPUs, resulting in zero cost.